# 1]organization=IIT-CNR, addressline=Via Moruzzi 1, code=56124, city=Pisa, country=Italy

### v3

@@ 7.2.5 Feature selection
We perform feature selection for all models except for the ensemble-based ones. Indeed, the latter perform internal feature selection, reducing the need for explicit preprocessing, and are robust to redundant features, unlike simpler models [69]. Feature selection is achieved with ANOVA F-value, a technique that was shown to achieve good performance in recent works on the detection of malicious online user activities [70]. The ANOVA tests are used to obtain an F-value for each feature, expressing the likelihood for the feature to be predictive for the task at hand. Then, features are ranked based on their respective F-value and the top-k are selected. Out of 142 total features, for each trained model we vary k from 10 to 80 with a step of 10, and evaluate the performance of the resulting models as part of the 10-fold cross-validation. At the end of the 10-fold cross-validation process, for each model we select the number and set of features that maximizes its F1 score on the positive class, so that for each model, only its best set of features is used in the following steps.
@@ 7.3 Model evaluation
Table 2: Classification results for the hard abandonment (top rows) and soft abandonment (bottom rows) tasks. Column p reports the number of features used by each model. Model performance is measured in terms of precision, recall and F1 on the positive class, and in terms of overall AUC and micro F1. For positive F1 and overall micro F1 we also report the 95% confidence intervals based on the scores obtained in the k-fold cross-validation. For each task, the best result in each evaluation metric is shown in bold and the second-best is underlined.
@@ 8.1 Prediction of hard and soft abandonment
Classification results for the hard abandonment task are reported in the top half of Table 2, while those for the soft abandonment task are reported in the bottom half of the table. For each task we report the performance of the baselines and of the optimized version of each trained model. For each baseline and model, Table 2 shows the number of used features (p) and the evaluation metrics computed for the positive class and for both classes (overall columns). With respect to the hard abandonment task, Gradient Boosting (GB) achieved the best results all-round. The second-best model in this task in terms of positive F1 is Random Forest (RF) that achieved competitive results in all metrics and second-best results in positive class recall and overall micro F1. Despite the relatively solid performance exhibited by some models (e.g., GB and RF), the results from Table 2 demonstrate the difficulty of this task. First of all, in general the results obtained by the trained models were encouraging but not exceptionally good, with the best result being micro F1 $=0.914$ by GB. Second, the difficulty faced by the trained models is also testified by the small margin by which some of the models —such as K-Nearest Neighbors (KNN) and Naive Bayes (NB)— outperformed the baselines, with Naive Bayes (NB) even having lower precision and positive F1 with respect to those achieved by DT-NFE in the hard abandonment task. However, for certain models our pipeline proved highly beneficial, significantly surpassing baseline performance as in the case of Support Vector Machine (SVM). Finally, results in table highlight a clear trend where the more complex models consistently outperformed the simpler ones. This shows that the increased complexity of models such as GB, RF, and AB was needed in order to obtain more accurate predictions. The scores reported in the bottom half of Table 2 reveal more nuanced results for the soft abandonment task. Again, the best-performing model is Gradient Boosting (GB), achieving positive F1 $=0.590$. Figure 5 shows precision-recall curves for this model, for both the hard and soft abandonment tasks. In hard abandonment, precision initially rises sharply and then stabilizes at a relatively high level for a wide range of recall values, before gradually decreasing. Conversely, in soft abandonment, while the initial trend is similar, precision declines more steeply as recall increases. This indicates that as the model attempts to retrieve more positive instances, it misclassifies a greater number of negative instances, leading to a faster deterioration in precision.
@@
A first interesting finding is that the overall results are worse for the soft abandonment task than for hard abandonment, while those for the positive class are better. For example, the best micro F1 $=0.814$ in soft abandonment, versus best micro F1 $=0.914$ in hard abandonment. On the contrary, for the simpler models that underwent our pipeline the best positive F1 $=0.499$ in soft abandonment, which is higher than F1 $=0.434$ in hard abandonment. In other words, by moving from hard to soft abandonment we improved predictions on the positive class (i.e., the abandoning users) but we degraded those on the negative one (i.e., the users who remained active). There are two related reasons for this behavior: (i) the definition of abandoning users and (ii) class imbalance. For the former, we recall that the soft abandonment task was introduced precisely to address the issue with those users who maintain some activity in the aftermath of the ban, but that eventually leave the platform, which represent challenging instances to classify. Therefore, their labeling in the soft abandonment task as abandoning users (positive class) rather than active ones (negative class), led to a certain degree of improvement of the models on the positive class. The second related factor is about class imbalance given that, before rebalancing, the soft abandonment task is less imbalanced than hard abandonment: 73/27 versus 85/15, respectively. To this end, we recall that abandoning users are the positive and the minority class in both tasks. As such, models trained for detecting soft abandonment train on a relatively larger number of positive instances than those for hard abandonment, hence the better performance on the positive class.
@@ 8.2 Predicting abandonment at different levels of user activity
We now delve deeper into the results by analyzing classification performance in relation to the level of user activity on the platform pre-ban. Evaluating classification performance based on user activity bears multiple implications, as it allows assessing whether classifiers exhibit varying levels of performance across different segments of users. For example, particularly active users —who exert the greatest influence on the platform— are pivotal in shaping community dynamics and behaviors [71, 50]. Therefore, correctly predicting their activity after an intervention provides moderators and platform administrators with valuable insights for their content management strategies. Conversely, accurately detecting the least active users —who are more likely to abandon the platform— enhances the utility of the prediction system by enabling proactive or alternative measures to retain users and mitigate churn [13].
@@
To perform this analysis, we examined the distribution of user comments pre-ban and we assign each user to one of the following five activity levels: very low (VL), low (LO), medium (ME), high (HI), and very high (VH) activity. The five levels are obtained by binning the distribution of user comments at regular intervals of 20 quantiles each, so that VL $\leq Q_{20}$, $Q_{20}<$ LO $\leq Q_{40}$, and so on up to VH $>Q_{80}$. We then performed a 80/20 train-test split of the dataset stratifying not only for class labels but also for the newly defined activity levels. Finally, we picked the best performing model based on the results in Table 2 —that is, Gradient Boosting (GB)— and we trained and optimized a different model for each activity level. The results of this analysis are reported in Table 3 for both the hard (top rows) and soft abandonment (bottom rows) task. In table are reported the number of comments n in each activity level, the corresponding number of abandoning (positive class) and remaining (negative class) users, and the evaluation metrics: the F1 score of the positive class, and the overall micro F1. The number of users in the positive and negative classes, reported in Table 3 for each activity level and for each task, shows that the imbalance in the dataset increases when considering increasingly active users. As we already observed in the results of Table 2, class imbalance in the data has a detrimental effect on the classification performance on the minority class (i.e., the positive class, corresponding to the abandoning users). In Table 3, this is reflected by the fact that the best positive F1 $=0.599$ is achieved on the soft abandonment task for users with very low (VL) activity, which is the experimental condition with the lowest class imbalance (61/39). However, this is not true for the hard abandonment task, in which the lowest positive F1 is obtained for those users whose activity level is low (LO), while the best are obtained for high (HI) and medium (ME) activity levels. The best overall result is achieved in the hard abandonment task for users with medium (ME) activity, with micro F1 $=0.941$. Then, in the soft abandonment task the best overall results are obtained at the medium (ME) and very high (VH) activity levels, with micro F1 $=0.842$ and $0.813$, respectively.
@@
hard abandon. VH HI --0.241 0.231 --0.311 0.580 0.280 0.420
HI ME --0.009 0.022 --0.329 0.323 0.258 0.389
ME LO --0.222 0.018 --0.263 0.296 0.245 0.375
LO VL --0.205 0.128 --0.415 0.239 0.235 0.340
VH VL --0.205 0.231 --0.291 0.135 0.250 0.295
HI LO --0.402 0.005 --0.329 --0.272 0.258 0.310
soft abandon. VH HI --0.352 0.128 --0.178 0.191 0.239 0.389
HI ME --0.188 0.005 --0.329 --0.014 0.267 0.357
ME LO --0.365 0.018 --0.263 --0.013 0.267 0.342
LO VL --0.120 0.128 --0.415 --0.038 0.235 0.330
VH VL --0.239 0.231 --0.291 0.191 0.250 0.307
HI LO --0.402 0.005 --0.329 --0.272 0.258 0.310
@@
After training a different model for each activity level, we assessed whether these models assigned similar importance to the features they used. We operationalized this analysis by computing several indicators of correlation and similarity between the feature rankings of the different models, such as Kendall’s $\tau$, Spearman’s $\rho$ and Rank-Biased Overlap (RBO) [72]. To rank the features of a model, we first identified and discarded strongly correlated features. In detail, we computed Pearson correlations r between each feature pair of a model, we identified those feature pairs for which $\textit{r}>|0.7|$, and we discarded one of such features at random. Then, for each model, we computed the feature importance of the remaining (uncorrelated) features with Permutation Importance (PI). PI measures the contribution of each feature to the model’s predictions by quantifying the performance degradation observed when its values are randomly shuffled while keeping all other features unchanged. Results are presented in Table 4 for the top-10 features of each model ($p=10$) and for all features ($p=\text{all}$). The column comparisons indicates which models —each corresponding to a different user activity level— are compared. As reported, we compared the feature ranking of models corresponding to similar activity levels (e.g., VH vs. HI), as well as those corresponding to opposite ones (e.g., VH vs. VL and HI vs. LO). The underlying intuition is that similar rankings should emerge among similar activity levels, whereas opposite levels are expected to exhibit less agreement.
@@
For the hard abandonment task, results in Table 4 show that considering all features led to higher agreement, suggesting that the correlation tended to increase towards the tail of the rankings rather than the head. For instance, the correlation between VH and HI showed strong agreement, with Spearman’s $\rho=0.580$ and RBO $=0.420$ when $p=\text{all}$, but decreased when $p=10$. This suggests that highly active users share similar characteristics and can be grouped together. For the adjacent levels, this agreement decreases, particularly when transitioning from HI to ME and further down the activity scale. For instance, HI and ME showed a noticeable drop in agreement, with Kendall’s $\tau=-0.009$, indicating that the ranking of features differs significantly. This suggests that while highly active users (VH and HI) may exhibit similar behaviors, comparing similar but lower ranges of activity levels introduces more variability. The agreement between adjacent levels, such as HI and ME or ME and LO, was notably weaker, especially in Kendall’s $\tau$, which remains close to zero or even negative. Surprisingly, the agreement between VH and VL is stronger than some adjacent-level comparisons, with Spearman’s $\rho=0.135$ and RBO $=0.295$. While this correlation is still moderate, it suggests that these two groups, despite their opposing activity levels, might share some behavioral patterns. A possible explanation is that both represent users with extreme behaviors, leading to similar feature relevance.
@@
Results in the soft abandonment task present the same patterns of the hard abandonment task. Higher agreement was generally found when all features are considered, with stronger correlations towards the tail of the rankings. For example, VH and HI exhibited relatively high agreement, particularly with respect to RBO, while Kendall’s $\tau$ remained lower, again highlighting sensitivity to rank shifts. Conversely, adjacent levels such as HI and ME, or ME and LO, show weak or even negative rank correlations, reinforcing the idea that mid-range activity users display more behavioral variability. Again, VH and VL show a correlation pattern similar to that of the hard abandonment task, with Spearman’s $\rho=0.191$ and RBO $=0.307$. This further supports the idea that extreme user groups, despite their contrasting activity levels, might share unexpected similarities. Overall, while some user groups exhibit similar feature importance rankings, others differ significantly. This highlights the need for tailored approaches rather than one-size-fits-all solutions. Understanding which users can be moderated similarly and which require separate treatment is pivotal for developing more personalized and effective moderation strategies [19].
@@ 8.3 Leave-one-out cross-validation
Until now, we trained our models on a dataset obtained by merging users from all 15 banned subreddits. However, users who participate in different subreddits could have different characteristics and exhibit different behaviors and reactions to a moderation intervention [14]. For this reason, we carried out an additional analysis aimed at assessing possible subreddit-specific factors influencing user abandonment. Importantly, this analysis also allows evaluating the extent to which a model trained on a certain set of subreddits is capable of generalizing to other —unseen— subreddits. To assess the generalizability of our best model we performed a leave-one-out cross-validation (LOOCV) by leveraging the multiplicity of banned subreddits in our dataset. This procedure represents the state-of-the-art in evaluating the generalizability of a classifier against multiple groups or classes of data instances [73], and has been recently used for related tasks such as the detection of different groups of bots [74, 75]. Based on the 15 banned subreddits in our dataset, we implemented the LOOCV by iteratively selecting 14 subreddits to train a Gradient Boosting (GB) hard and soft abandonment detection classifier, and by testing the trained model on users from the remaining (held-out) subreddit. For this analysis, we considered a user to participate in a subreddit if it posted more than 10 comments in that subreddit, so as to reduce noise caused by sporadic participation.
@@
The validity of the LOOCV hinges on the assumption of independence between the training and testing datasets. Specifically, to ensure unbiased evaluation, it is required that the held-out subreddit contains users who are largely distinct from those in the training subreddits. While the independence assumption is easily verified in some domains of application, such as that of bot detection where each bot belongs to only one botnet [74], it does not hold in our context where any user can participate in multiple subreddits. When a considerable overlap exists between the sets of users who participate in two subreddits, utilizing either of these subreddits as the held-out test dataset may leak information to the classifier from the training datasets, thereby compromising the validity of the evaluation [76]. Before running the LOOCV we thus assessed the extent of overlap between the sets of users who participated in the 15 banned subreddits. The results of this analysis are presented in Figure 6 as a heatmap of the user overlap matrix between the subreddits. In figure, each cell $o_{ij}$ in the user overlap matrix $O$ reports the percentage of users from the $i$-th subreddit that also participated in the $j$-th subreddit. By definition, cells in the matrix diagonal correspond to 100% overlap and are dark-colored, while those that correspond to no overlap are white colored. Apart from r/ShitNeoconsSay (SNS) and r/soyboys (SB) that share up to 75% of their users with r/ConsumeProducts (CP), the matrix is overall very sparse, as demonstrated by the limited reported overlap. The results from this analysis show that the sets of users who participated in the banned subreddits are largely disjoint, in line with the literature on echo chambers [77], which supports the application of the LOOCV.
@@
Table 5 shows the results of the LOOCV for the hard (top rows) and soft abandonment (bottom rows) tasks. Each row in the table reports testing results on a specific subreddit, when training on all others. In addition to the evaluation metrics, for each subreddit we also report the number of participating users and its maximum overlap with the other subreddits. Finally, for each task, the bottom rows report evaluation results aggregated across all subreddits, in terms of mean and standard deviation of each metric, as well the scores obtained by the GB model trained and tested on data from all subreddits —that is, without the LOOCV— as reported in Table 2. The latter scores serve as reference values with which to compare the LOOCV results to quantify the performance decrease to expect when classifying users from an unseen subreddit. The averaged results reported in Table 5 for both tasks indicate a high variability in the evaluation metrics of the positive class, as demonstrated by the large standard deviation with respect to the means. For instance, in the hard abandonment task, the F1 score ranges from a minimum of $0.000$ for r/TheNewRight (TNR) and r/imgoingtohellforthis2 (ITH2) to a maximum of $0.846$ for r/Wojak (WJ). Similarly, the recall ranges from $0.000$ for r/TheNewRight (TNR) and r/imgoingtohellforthis2 (ITH2) to $1.000$ for r/OandAExclusiveForum (OAE). The same can be observed for the soft abandonment task as well, where the precision ranges from $0.184$ for r/ChapoTrapHouse (CTH) to $0.827$ for r/GenderCritical (GC). These results are indicative of substantial differences among the banned subreddits. In fact, in addition to the overall limited user overlap observed in Figure 6, the results from Table 5 highlight notable differences in the behavior exhibited by the participants of such subreddits. Interestingly enough, the maximum overlap that each subreddit has with others, has little influence on the performance of the trained model. For each task, we computed the Pearson correlation coefficient r to estimate the strength of the linear relationship between the max overlap and two evaluation metrics —namely, the F1 score of the positive class and the micro F1. The largest value measured for the hard abandonment task is r $=0.237$ (p $=0.394$), while the largest measured for the soft abandonment task is r $=0.222$ (p $=0.493$), both of which indicate weak and non-significant correlations. In turn, this strengthens the results reported in Table 5, including those about subreddit diversity, as the variable results in the table cannot be simply explained by the fact that some users participated in multiple subreddits, but rather to their inherent differences. The comparison between the results obtained without LOOCV and the LOOCV aggregate results sheds light on the capacity of our model to generalize to unseen data. To this end, results in Table 5 show a moderate loss in performance in the LOOCV experiment, which is expected given the aforementioned differences among the subreddits. However, the percentage loss is generally contained. For example, in the hard abandonment task the average F1 score on the positive class obtained with the LOOCV is 60.0% of the original one. The loss for the overall metrics is even lower, with the average micro F1 reaching 85.0% of the original value. Similar percentage losses are obtained for the soft abandonment task in the overall metrics, with the LOOCV average positive F1 maintaining 90.7% of the original value. Overall, these figures suggest that the model maintains encouraging performance even when applied to unseen data.

### v4

@@ 7.2.5 Feature selection
We perform feature selection for all models except for the ensemble-based ones. Indeed, the latter perform internal feature selection, reducing the need for explicit preprocessing, and are robust to redundant features, unlike simpler models [bolon2019ensembles]. Feature selection is achieved with ANOVA F-value, a technique that was shown to achieve good performance in recent works on the detection of malicious online user activities [ilias2021detecting]. The ANOVA tests are used to obtain an F-value for each feature, expressing the likelihood for the feature to be predictive for the task at hand. Then, features are ranked based on their respective F-value and the top-k are selected. Out of 142 total features, for each trained model we vary k from 10 to 80 with a step of 10, and evaluate the performance of the resulting models as part of the 10-fold cross-validation. At the end of the 10-fold cross-validation process, for each model we select the number and set of features that maximizes its F1 score on the positive class, so that for each model, only its best set of features is used in the following steps.
@@ 7.3 Model evaluation
Table 2: Classification results for the hard abandonment (top rows) and soft abandonment (bottom rows) tasks. Column p reports the number of features used by each model. Model performance is measured in terms of precision, recall and F1 on the positive class, and in terms of overall AUC and micro F1. For positive F1 and overall micro F1 we also report the 95% confidence intervals based on the scores obtained in the k-fold cross-validation.
For each task, the best result in each evaluation metric is shown in bold and the second-best is underlined.
@@ 8.1 Prediction of hard and soft abandonment
Classification results for the hard abandonment task are reported in the top half of Table 2, while those for the soft abandonment task are reported in the bottom half of the table. For each task we report the performance of the baselines and of the optimized version of each trained model. For each baseline and model, Table 2 shows the number of used features (p) and the evaluation metrics computed for the positive class and for both classes (overall columns). With respect to the hard abandonment task, Gradient Boosting (GB) achieved the best results all-round. The second-best model in this task in terms of positive F1 is Random Forest (RF) that achieved competitive results in all metrics and second-best results in positive class recall and overall micro F1. Despite the relatively solid performance exhibited by some models (e.g., GB and RF), the results from Table 2 demonstrate the difficulty of this task. First of all, in general the results obtained by the trained models were encouraging but not exceptionally good, with the best result being micro F1 =0.914=0.914 by GB. Second, the difficulty faced by the trained models is also testified by the small margin by which some of the models —such as K-Nearest Neighbors (KNN) and Naive Bayes (NB)— outperformed the baselines, with Naive Bayes (NB) even having lower precision and positive F1 with respect to those achieved by DT-NFE in the hard abandonment task. However, for certain models our pipeline proved highly beneficial, significantly surpassing baseline performance as in the case of Support Vector Machine (SVM). Finally, results in table highlight a clear trend where the more complex models consistently outperformed the simpler ones. This shows that the increased complexity of models such as GB, RF, and AB was needed in order to obtain more accurate predictions. The scores reported in the bottom half of Table 2 reveal more nuanced results for the soft abandonment task. Again, the best-performing model is Gradient Boosting (GB), achieving positive F1 =0.590=0.590. Figure 5 shows precision-recall curves for this model, for both the hard and soft abandonment tasks. In hard abandonment, precision initially rises sharply and then stabilizes at a relatively high level for a wide range of recall values, before gradually decreasing. Conversely, in soft abandonment, while the initial trend is similar, precision declines more steeply as recall increases. This indicates that as the model attempts to retrieve more positive instances, it misclassifies a greater number of negative instances, leading to a faster deterioration in precision.
A first interesting finding is that the overall results are worse for the soft abandonment task than for hard abandonment, while those for the positive class are better. For example, the best micro F1 =0.814=0.814 in soft abandonment, versus best micro F1 =0.914=0.914 in hard abandonment.
@@
On the contrary, for the simpler models that underwent our pipeline the best positive F1 =0.499=0.499 in soft abandonment, which is higher than F1 =0.434=0.434 in hard abandonment. In other words, by moving from hard to soft abandonment we improved predictions on the positive class (i.e., the abandoning users) but we degraded those on the negative one (i.e., the users who remained active). There are two related reasons for this behavior: (i) the definition of abandoning users and (ii) class imbalance. For the former, we recall that the soft abandonment task was introduced precisely to address the issue with those users who maintain some activity in the aftermath of the ban, but that eventually leave the platform, which represent challenging instances to classify. Therefore, their labeling in the soft abandonment task as abandoning users (positive class) rather than active ones (negative class), led to a certain degree of improvement of the models on the positive class. The second related factor is about class imbalance given that, before rebalancing, the soft abandonment task is less imbalanced than hard abandonment: 73/27 versus 85/15, respectively. To this end, we recall that abandoning users are the positive and the minority class in both tasks. As such, models trained for detecting soft abandonment train on a relatively larger number of positive instances than those for hard abandonment, hence the better performance on the positive class.
@@ 8.2 Predicting abandonment at different levels of user activity
We now delve deeper into the results by analyzing classification performance in relation to the level of user activity on the platform pre-ban. Evaluating classification performance based on user activity bears multiple implications, as it allows assessing whether classifiers exhibit varying levels of performance across different segments of users. For example, particularly active users —who exert the greatest influence on the platform— are pivotal in shaping community dynamics and behaviors [zignani2019mastodon, robertson2022uncommon]. Therefore, correctly predicting their activity after an intervention provides moderators and platform administrators with valuable insights for their content management strategies. Conversely, accurately detecting the least active users —who are more likely to abandon the platform— enhances the utility of the prediction system by enabling proactive or alternative measures to retain users and mitigate churn [trujillo2022make].
@@
To perform this analysis, we examined the distribution of user comments pre-ban and we assign each user to one of the following five activity levels: very low (VL), low (LO), medium (ME), high (HI), and very high (VH) activity. The five levels are obtained by binning the distribution of user comments at regular intervals of 20 quantiles each, so that VL $\leq Q_{20}$, $Q_{20}<$ LO $\leq Q_{40}$, and so on up to VH $Q_{80}$. We then performed a 80/20 train-test split of the dataset stratifying not only for class labels but also for the newly defined activity levels. Finally, we picked the best performing model based on the results in Table 2 —that is, Gradient Boosting (GB)— and we trained and optimized a different model for each activity level. The results of this analysis are reported in Table 3 for both the hard (top rows) and soft abandonment (bottom rows) task. In table are reported the number of comments n in each activity level, the corresponding number of abandoning (positive class) and remaining (negative class) users, and the evaluation metrics: the F1 score of the positive class, and the overall micro F1. The number of users in the positive and negative classes, reported in Table 3 for each activity level and for each task, shows that the imbalance in the dataset increases when considering increasingly active users. As we already observed in the results of Table 2, class imbalance in the data has a detrimental effect on the classification performance on the minority class (i.e., the positive class, corresponding to the abandoning users). In Table 3, this is reflected by the fact that the best positive F1 =0.599=0.599 is achieved on the soft abandonment task for users with very low (VL) activity, which is the experimental condition with the lowest class imbalance (61/39). However, this is not true for the hard abandonment task, in which the lowest positive F1 is obtained for those users whose activity level is low (LO), while the best are obtained for high (HI) and medium (ME) activity levels. The best overall result is achieved in the hard abandonment task for users with medium (ME) activity, with micro F1 =0.941=0.941. Then, in the soft abandonment task the best overall results are obtained at the medium (ME) and very high (VH) activity levels, with micro F1 =0.842=0.842 and $0.813$, respectively.
@@
hard abandon. VH HI -0.241 0.231 -0.311 0.580 0.280 0.420
HI ME -0.009 0.022 -0.329 0.323 0.258 0.389
ME LO -0.222 0.018 -0.263 0.296 0.245 0.375
LO VL -0.205 0.128 -0.415 0.239 0.235 0.340
VH VL -0.205 0.231 -0.291 0.135 0.250 0.295
HI LO -0.402 0.005 -0.329 -0.272 0.258 0.310
soft abandon. VH HI -0.352 0.128 -0.178 0.191 0.239 0.389
HI ME -0.188 0.005 -0.329 -0.014 0.267 0.357
ME LO -0.365 0.018 -0.263 -0.013 0.267 0.342
LO VL -0.120 0.128 -0.415 -0.038 0.235 0.330
VH VL -0.239 0.231 -0.291 0.191 0.250 0.307
HI LO -0.402 0.005 -0.329 -0.272 0.258 0.310
@@
After training a different model for each activity level, we assessed whether these models assigned similar importance to the features they used. We operationalized this analysis by computing several indicators of correlation and similarity between the feature rankings of the different models, such as Kendall’s $\tau$, Spearman’s $\rho$ and Rank-Biased Overlap (RBO) [webber2010similarity]. To rank the features of a model, we first identified and discarded strongly correlated features. In detail, we computed Pearson correlations r between each feature pair of a model, we identified those feature pairs for which $\textit{r}>|0.7|$, and we discarded one of such features at random. Then, for each model, we computed the feature importance of the remaining (uncorrelated) features with Permutation Importance (PI). PI measures the contribution of each feature to the model’s predictions by quantifying the performance degradation observed when its values are randomly shuffled while keeping all other features unchanged. Results are presented in Table 4 for the top-10 features of each model ($p=10$) and for all features ($p=\text{all}$). The column comparisons indicates which models —each corresponding to a different user activity level— are compared. As reported, we compared the feature ranking of models corresponding to similar activity levels (e.g., VH vs. HI), as well as those corresponding to opposite ones (e.g., VH vs. VL and HI vs. LO). The underlying intuition is that similar rankings should emerge among similar activity levels, whereas opposite levels are expected to exhibit less agreement.
@@
For the hard abandonment task, results in Table 4 show that considering all features led to higher agreement, suggesting that the correlation tended to increase towards the tail of the rankings rather than the head. For instance, the correlation between VH and HI showed strong agreement, with Spearman’s $\rho=0.580$ and RBO =0.420=0.420 when $p=\text{all}$, but decreased when $p=10$. This suggests that highly active users share similar characteristics and can be grouped together. For the adjacent levels, this agreement decreases, particularly when transitioning from HI to ME and further down the activity scale. For instance, HI and ME showed a noticeable drop in agreement, with Kendall’s $\tau=-0.009$, indicating that the ranking of features differs significantly. This suggests that while highly active users (VH and HI) may exhibit similar behaviors, comparing similar but lower ranges of activity levels introduces more variability. The agreement between adjacent levels, such as HI and ME or ME and LO, was notably weaker, especially in Kendall’s $\tau$, which remains close to zero or even negative. Surprisingly, the agreement between VH and VL is stronger than some adjacent-level comparisons, with Spearman’s $\rho=0.135$ and RBO =0.295=0.295. While this correlation is still moderate, it suggests that these two groups, despite their opposing activity levels, might share some behavioral patterns. A possible explanation is that both represent users with extreme behaviors, leading to similar feature relevance.
@@
Results in the soft abandonment task present the same patterns of the hard abandonment task. Higher agreement was generally found when all features are considered, with stronger correlations towards the tail of the rankings. For example, VH and HI exhibited relatively high agreement, particularly with respect to RBO, while Kendall’s $\tau$ remained lower, again highlighting sensitivity to rank shifts. Conversely, adjacent levels such as HI and ME, or ME and LO, show weak or even negative rank correlations, reinforcing the idea that mid-range activity users display more behavioral variability. Again, VH and VL show a correlation pattern similar to that of the hard abandonment task, with Spearman’s $\rho=0.191$ and RBO =0.307=0.307. This further supports the idea that extreme user groups, despite their contrasting activity levels, might share unexpected similarities. Overall, while some user groups exhibit similar feature importance rankings, others differ significantly. This highlights the need for tailored approaches rather than one-size-fits-all solutions. Understanding which users can be moderated similarly and which require separate treatment is pivotal for developing more personalized and effective moderation strategies [cresci2022personalized].
@@ 8.3 Leave-one-out cross-validation
Until now, we trained our models on a dataset obtained by merging users from all 15 banned subreddits. However, users who participate in different subreddits could have different characteristics and exhibit different behaviors and reactions to a moderation intervention [trujillo2023one]. For this reason, we carried out an additional analysis aimed at assessing possible subreddit-specific factors influencing user abandonment. Importantly, this analysis also allows evaluating the extent to which a model trained on a certain set of subreddits is capable of generalizing to other —unseen— subreddits. To assess the generalizability of our best model we performed a leave-one-out cross-validation (LOOCV) by leveraging the multiplicity of banned subreddits in our dataset. This procedure represents the state-of-the-art in evaluating the generalizability of a classifier against multiple groups or classes of data instances [wong2015performance], and has been recently used for related tasks such as the detection of different groups of bots [echeverri2018lobo, mannocci2022mulbot]. Based on the 15 banned subreddits in our dataset, we implemented the LOOCV by iteratively selecting 14 subreddits to train a Gradient Boosting (GB) hard and soft abandonment detection classifier, and by testing the trained model on users from the remaining (held-out) subreddit. For this analysis, we considered a user to participate in a subreddit if it posted more than 10 comments in that subreddit, so as to reduce noise caused by sporadic participation.
@@
The validity of the LOOCV hinges on the assumption of independence between the training and testing datasets. Specifically, to ensure unbiased evaluation, it is required that the held-out subreddit contains users who are largely distinct from those in the training subreddits. While the independence assumption is easily verified in some domains of application, such as that of bot detection where each bot belongs to only one botnet [echeverri2018lobo], it does not hold in our context where any user can participate in multiple subreddits. When a considerable overlap exists between the sets of users who participate in two subreddits, utilizing either of these subreddits as the held-out test dataset may leak information to the classifier from the training datasets, thereby compromising the validity of the evaluation [cresci2024demystifying]. Before running the LOOCV we thus assessed the extent of overlap between the sets of users who participated in the 15 banned subreddits. The results of this analysis are presented in Figure 6 as a heatmap of the user overlap matrix between the subreddits. In figure, each cell $o_{ij}$ in the user overlap matrix $O$ reports the percentage of users from the $i$-th subreddit that also participated in the $j$-th subreddit. By definition, cells in the matrix diagonal correspond to 100% overlap and are dark-colored, while those that correspond to no overlap are white colored. Apart from r/ShitNeoconsSay (SNS) and r/soyboys (SB) that share up to 75% of their users with r/ConsumeProducts (CP), the matrix is overall very sparse, as demonstrated by the limited reported overlap. The results from this analysis show that the sets of users who participated in the banned subreddits are largely disjoint, in line with the literature on echo chambers [cinelli2021echo], which supports the application of the LOOCV.
@@
Table 5 shows the results of the LOOCV for the hard (top rows) and soft abandonment (bottom rows) tasks. Each row in the table reports testing results on a specific subreddit, when training on all others. In addition to the evaluation metrics, for each subreddit we also report the number of participating users and its maximum overlap with the other subreddits. Finally, for each task, the bottom rows report evaluation results aggregated across all subreddits, in terms of mean and standard deviation of each metric, as well the scores obtained by the GB model trained and tested on data from all subreddits —that is, without the LOOCV— as reported in Table 2. The latter scores serve as reference values with which to compare the LOOCV results to quantify the performance decrease to expect when classifying users from an unseen subreddit. The averaged results reported in Table 5 for both tasks indicate a high variability in the evaluation metrics of the positive class, as demonstrated by the large standard deviation with respect to the means. For instance, in the hard abandonment task, the F1 score ranges from a minimum of $0.000$ for r/TheNewRight (TNR) and r/imgoingtohellforthis2 (ITH2) to a maximum of $0.846$ for r/Wojak (WJ). Similarly, the recall ranges from $0.000$ for r/TheNewRight (TNR) and r/imgoingtohellforthis2 (ITH2) to $1.000$ for r/OandAExclusiveForum (OAE). The same can be observed for the soft abandonment task as well, where the precision ranges from $0.184$ for r/ChapoTrapHouse (CTH) to $0.827$ for r/GenderCritical (GC). These results are indicative of substantial differences among the banned subreddits. In fact, in addition to the overall limited user overlap observed in Figure 6, the results from Table 5 highlight notable differences in the behavior exhibited by the participants of such subreddits. Interestingly enough, the maximum overlap that each subreddit has with others, has little influence on the performance of the trained model. For each task, we computed the Pearson correlation coefficient r to estimate the strength of the linear relationship between the max overlap and two evaluation metrics —namely, the F1 score of the positive class and the micro F1. The largest value measured for the hard abandonment task is r =0.237=0.237 (p =0.394=0.394), while the largest measured for the soft abandonment task is r =0.222=0.222 (p =0.493=0.493), both of which indicate weak and non-significant correlations. In turn, this strengthens the results reported in Table 5, including those about subreddit diversity, as the variable results in the table cannot be simply explained by the fact that some users participated in multiple subreddits, but rather to their inherent differences. The comparison between the results obtained without LOOCV and the LOOCV aggregate results sheds light on the capacity of our model to generalize to unseen data. To this end, results in Table 5 show a moderate loss in performance in the LOOCV experiment, which is expected given the aforementioned differences among the subreddits. However, the percentage loss is generally contained. For example, in the hard abandonment task the average F1 score on the positive class obtained with the LOOCV is 60.0% of the original one. The loss for the overall metrics is even lower, with the average micro F1 reaching 85.0% of the original value. Similar percentage losses are obtained for the soft abandonment task in the overall metrics, with the LOOCV average positive F1 maintaining 90.7% of the original value. Overall, these figures suggest that the model maintains encouraging performance even when applied to unseen data.